suppose both a 6% and 12% alcohol solution are available. How many gallons of each must be mixed in order to get 30 gallons of a 10.5% alcohol solution?
Answer by richwmiller(17219)

We need a total of 30 gal at 10.5%
We want to know how many gal at 6% to mix with 12%
a+b=30,
0.06*a+0.12*b=0.105*30
a=30-b
0.06*(30-b)+0.12*b=3.15
1.8-0.06b+0.12*b=3.15
0.06*b=1.35
b=22.5
a=30-b
a=7.5 gal at 6%
b=22.5 gal at 12%
check
0.06*7.5+0.12*22.5=0.105*30
0.45+2.7=3.15
3.15=3.15
ok
codemx1